OpenStudy (anonymous):

Graph the system of linear inequalities below. 3x+y≤ 6 4x-2y>6

OpenStudy (anonymous):

What do you think ?

OpenStudy (anonymous):

im not sure:/

OpenStudy (anonymous):

This will help :-) Let me know if you have any questions. https://www.desmos.com/calculator/nbdcxgdjta

OpenStudy (anonymous):

I plotted it for you.

OpenStudy (anonymous):

ok do i copy it on my graph then

OpenStudy (anonymous):

Yes, you can do that.

OpenStudy (anonymous):

What I gave you is what the graph is going to look like. The area that over laps are the solutions.

OpenStudy (anonymous):

so (1.5,0) and (0,-3)

OpenStudy (anonymous):

1.8 and 0.6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What are those points for??

OpenStudy (anonymous):

its from the graph u gave me

OpenStudy (anonymous):

do i need points for this equation?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

To graph it yes on paper yes. TO do that you need to turn the equilities into equations and then plot the lines For instance 3x+y≤ 6 = 3x+y=6 4x-2y>6 = 4x-2y= 6 Now we plot these equation of a line 3x+y=6 4x-2y= 6 Lets do the first equation of the line 3x+y=6 When x = 1 3x+y=6 3(1)+y=6 3(1)+y=6 3+y=6 y=6-3 y=3 So we have are first point for that line/equation which is (1,3) but we need one or two more to draw our line. Use the number 2 instead of one but do it the same way I did for number 1. Do you understand?

OpenStudy (anonymous):

You can also use your mouse on that site I gave you and click on the line and hold down to drag and you can see the points

OpenStudy (anonymous):

and the 4x-2y= 6 would be 0,3 right @Nixy